OpenStudy (anonymous):

The twice-differentiable function f is defined for all real numbers and satisfies the following conditions: f(0)= 4, f'(0)= -3, and f''(0)= 2 a) the function is given by g(x)= a ln(f(x)) for all real numbers, where a is a constant.Find g'(0) and g''(0) in terms of a. Show the work that leads to your answers b) the function h is given by h(x) =sec(kx)f(x) for all real numbers, where k is a constant. Find h'(x) and write an equation for the line tangent to the graph of h at x=0

OpenStudy (alekos):

to start off with g'(x) = af'(x)/f(x) so g(0) = -3a/4

OpenStudy (alekos):

sorry g'(0) = -3a/4

OpenStudy (raffle_snaffle):

Do you know how to differentiate?

OpenStudy (alekos):

my working goes like this......... g(x)= a ln[f(x)] the derivative of lnx is 1/x so the derivative of ln[f(x)] should be [1/f(x)].f'(x) hence g'(x) = af'(x)/f(x)

OpenStudy (alekos):

let me know if you think this is wrong

OpenStudy (alekos):

g''(x) is a little more complex

OpenStudy (alekos):

I have an answer for this as well but i'll let dmendez carry on from here

OpenStudy (anonymous):

Thanks for your help, I didn't understand what the a in the first given function meant but now I know how to find the derivative of F(x) to find g'(x)

OpenStudy (anonymous):

I'm having trouble finding g"(x) though

OpenStudy (alekos):

yeah its tricky you have to use the quotient rule g'(u/v) = (u'v-v'u)/v^2 where u = af'(x) and v = f(x)

OpenStudy (alekos):

you should get g''(0) = -a/16

OpenStudy (alekos):

give it a try and see how you go while i work on the second part

OpenStudy (anonymous):

okay, thanks

OpenStudy (anonymous):

okay so, the derivative of g(x) would be ((a f'(x))/(f(x))

OpenStudy (alekos):

yes, that's right, but what did you get for the 2nd derivative of g(x) ?

OpenStudy (anonymous):

yeah, would it be (a f"(x)=(a f'(x)^2)/(f(x))+f(x) g"(x))?

OpenStudy (alekos):

no its........ g'' = (af''f -af'^2)/f^2 i've left out the x's for simplicity

OpenStudy (anonymous):

is (a(-f'(x)^2)+f(x)f"(x)))/f(x)^2=g"(x)) correct?

OpenStudy (alekos):

yes thats it!

OpenStudy (anonymous):

Cool, Thank you so much for your help!

OpenStudy (alekos):

no problem. i'll let you do the rest. if you need help with the 2nd part let me know

OpenStudy (anonymous):

thanks, the second part I understand:)